# Evidence for simple “arrow of time functions” in closed chaotic quantum systems

### v1

@@ IV.1 The Recursion Method and Lanczos coefficients
where $b_{0}=0$. Upon defining the functions
@@
the connection to the scheme (1) becomes obvious. The time derivatives of the autocorrelation function in the AOTF construction correspond to the repeated action of the Liouvillian onto the respective observable. Thus the functions $\varphi_{n}(t)$ are equal to the overlap of the $n^{\text{th}}$ Krylov vector with $\mathcal{O}(t)$ as explained in (21). The quantities $b_{n}$ are by construction real and positive and are referred to as Lanczos coefficients.
With the definition (21) comes a frequently adapted interpretation of the Lanczos coefficients as hopping amplitudes of a tight-binding model. The entries of the wave function on this semi-infinite chain are given by (21). For a sketch see Fig. 10. The wave vector satisfies a discrete Schrödinger equation
@@
where we set $\varphi_{-1}=0$ by convention and with initial condition $\varphi_{n}(0)=\delta_{n,0}$. This may be cast into the form of a matrix equation
@@ V.1 Replacement model
Correspondingly, for the respective quantity in the Lanczos picture we compute
@@
where $\textbf{n}=\text{diag}(1,2,\dots)$, again suggesting the form of the potential to be $V(x)=-b^{2}(x)$.
@@
In order for a particle to propagate and to not be trapped, the potential $V$ necessarily needs to fall off, i.e. the Lanczos coefficients need to ascend. Furthermore, on qualitative grounds we argue that for a steady “forward-oriented” motion of the particle, in the sense of no back-transport, the potential additionally needs to exhibit a certain smoothness in order to not give rise to scattering or localization. For this second property we deliberately remain unspecific as we only aim to further stress the suitability of the replacement model rather than to quantify notions of smoothness.
@@ V.2 Operator Growth Hypothesis
For numerical purposes however, one is limited to finite systems. In such systems of size $L$, the first $L/2$ Lanczos coefficients $b_{n}$ coincide with those of the infinite system. Several studies suggest that already the first several Lanczos coefficients feature a linear trend [16, 28, 29, 30, 31]. However whether this precise behaviour persists forever clearly lies beyond the scope of any numerical investigation and is no part of the OGH.
@@ V.3 Lanczos Coefficients and Intricacies
In this work we proposed a construction for monotonously decreasing arrow of time functions (AOTFs) for autocorrelation functions of local observables. The construction of the AOTFs requires $2\mathcal{I}$ temporal derivatives of the respective autocorrelation function itself. Hence the intricacy $\mathcal{I}$ characterises the extend to which an AOTF may be considered to be time-local with respect to the autocorrelation function.
We numerically investigated the AOTFs of several few-body observables in pertinent quantum models. In these we found that whenever the system is fully chaotic (subject to the gap ratio) the intricacies $\mathcal{I}$ are low, allowing for a time-local description of the systems’ approach to equilibrium. Conversely, only when the system is nonchaotic or close to a nonchaotic regime (with respect to some model parameter) the dynamics may become highly intricate. However, integrability itself (or closeness to it) is no sufficient condition for high-intricacy dynamics.
@@ Acknowledgements
We thank Christian Bartsch, Robin Heveling and Mats H. Lamann for fruitful discussions. JW thanks Heiko G. Menzler for discussions on the numerical method. This work has been funded by the Deutsche Forschungsgemeinschaft (DFG), under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190.
Appendix A Numerical method

### v3

@@ IV.1 The Recursion Method and Lanczos coefficients
where $b_{0}=0$. The quantities $b_{n}$ are by construction real and positive and are referred to as Lanczos coefficients. Upon defining the functions
@@
we may highlight the connection of the Lanczos algorithm defined Eq. (20) and the construction of the AOTFs in Eq. (1). The time derivatives of the autocorrelation function in the AOTF construction are related to the repeated action of the Liouvillian onto the respective observable. Starting from the Heisenberg equation of motion for $\mathcal{O}(t)$ it may be shown (cf. e.g., Refs. [17, 38]) that the “wave vector” satisfies a discrete Schrödinger equation
@@
where we set $\varphi_{-1}=0$ by convention and with initial condition $\varphi_{n}(0)=\delta_{n,0}$. Note that Eq. (22) is equivalent to the construction given in (1). Further technical details on the equivalence between the algorithms described in Eq. (1) and Eqs. (20),(21) are provided in Appendix A. The functions $\varphi_{n}(t)$ are equal to the overlap of the $n^{\text{th}}$ Krylov vector with $\mathcal{O}(t)$ as explained in (21). For an example of their (wavelike) dynamics see Fig. 10.
For later reference we note that this may be cast into the form of a matrix equation
@@
Renouncing on the “gauge” (21) yields, as an analogue of (22),
$i\partial_{t}(\mathcal{O}_{n}|\mathcal{O}(t))=-\left[b_{n}(\mathcal{O}_{n-1}|\mathcal{O}(t))+b_{n+1}(\mathcal{O}_{n+1}|\mathcal{O}(t))\right].$
This (other) discrete Schrödinger equation suggests a frequently adapted interpretation of the Lanczos coefficients as hopping amplitudes of a tight-binding model, cf. Fig  11. The components of the wave function for the $n$-th site on this semi-infinite chain are given by the $(\mathcal{O}_{n}|\mathcal{O}(t))$. As $\varphi_{n}^{2}(t)=|(\mathcal{O}_{n}|\mathcal{O}(t))|^{2}$ holds, the dynamics of the $\varphi_{n}(t)$ as generated by (22) are equivalent to those of a particle hopping on a chain as long as only the properties to sit on specific sites are addressed.
@@ IV.2 Currents and probabilities
Furthermore the AOTF’s are related to the concept of Krylov-complexity $K$. With the notation as above the Krylov-complexity reads
$K(t)=\sum_{n=0}^{\infty}n\ \boldsymbol{\varphi}_{n}^{2}(t),$ (29)
which in this setting may be interpreted as the expectation value of an observable $\textbf{N}=\text{diag}(0,1,2,\dots)$, i.e. $\langle\boldsymbol{\varphi}(t)|\textbf{N}|\boldsymbol{\varphi}(t)\rangle=K(t)$. Again invoking the analogy of the Lanczos picture to a particle hopping along a semi-infinite lead as visualized in Fig. 11, we may also interpret $K(t)$ as the expectation value of the position of the particle. Within this framework the functions $\alpha^{n}(t)=\langle\textbf{P}_{n}(t)\rangle$ from (2) and (26), respectively, may be interpreted as the probability to find the particle at any position at or left of $n$  [^4] In fact $\alpha^{\mathcal{I}}(t)$ itself is a $q$-complexity as $\alpha^{\mathcal{I}}(t)=\sum_{n=0}^{\infty}\Theta(\mathcal{I}-n)\boldsymbol{\varphi}^{2}_{n}(t)$..
@@ V.1 Replacement model
Figure 11: Illustration of the replacement model for the Lanczos dynamics. In the semi-infinite tight-binding model (left) the Lanczos coefficients $b_{n}$ may be viewed as discrete hopping amplitudes. The probability of the particle to be located at some site $j$ is given by $\varphi^{2}_{j}$. In the replacement model (see sketch on the right) a continuous version of the Lanczos coefficients takes on the role of a potential in the form of $V=-b^{2}(x)$. This picture motivates the necessity of growing Lanczos coefficients (i.e. a decreasing potential) in order for a particle with mass $m=1/2$ to propagate outbound. This connection relates general properties of the respective Lanczos coefficients with the dynamics of the wave vector $\boldsymbol{\varphi}(t)$ and hence to the arrow of time functions $\alpha^{\mathcal{I}}$ via (27).
@@
To further strengthen this analogy we consider the dynamics of $\langle\hat{x}\rangle$ in the replacement model. By the Ehrenfest theorem we have for a particle with mass $m=1/2$ that
@@
Correspondingly, for the respective quantity in the Lanczos picture, i.e. , the Krylov-complexity, we compute
@@
with N as defined below Eq. 29, again suggesting the form of the potential to be $V(x)=-b^{2}(x)$.
@@
In order for a particle to propagate and to not be trapped, the potential $V$ necessarily needs to fall off, i.e. the Lanczos coefficients need to ascend. Qualitatively similar statements may be found in Ref. [17]. Furthermore, on qualitative grounds we argue that for a steady “forward-oriented” motion of the particle, in the sense of no back-transport, the potential additionally needs to exhibit a certain smoothness in order to not give rise to scattering or localization. For this second property we deliberately remain unspecific as we only aim to further stress the suitability of the replacement model rather than to quantify notions of smoothness.
@@ V.2 Operator Growth Hypothesis
For numerical purposes however, one is limited to finite systems. In such systems of size $L$, the first $L/2$ Lanczos coefficients $b_{n}$ coincide with those of the infinite system. Several studies suggest that already the first several Lanczos coefficients feature a linear trend [17, 42, 43, 38, 44].
@@ V.3 Lanczos Coefficients and Intricacies
In this work we proposed a construction for monotonously decreasing arrow of time functions (AOTFs) for autocorrelation functions of local observables. The construction of the AOTFs requires $2\mathcal{I}$ temporal derivatives of the respective autocorrelation function itself. Due to this, and our qualitative assessment that autocorrelation function yielding larger $\mathcal{I}$ feature richer structures, we refer to $\mathcal{I}$ as intricacy.
We numerically investigated the AOTFs of several few-body observables in pertinent quantum models. In these we found that whenever the system is fully chaotic (subject to the gap ratio) the intricacies $\mathcal{I}$ are low. Conversely, only when the system is nonchaotic or close to a nonchaotic regime (with respect to some model parameter) the dynamics may become highly intricate. However, integrability itself (or closeness to it) is no sufficient condition for high-intricacy dynamics.
@@ Acknowledgements
We thank Christian Bartsch, Robin Heveling, Mats H. Lamann, Alvaro M. Alhambra, Robin Steinigeweg and Philipp Strasberg for fruitful discussions. J.W. thanks Heiko G. Menzler for discussions on the numerical method. This work has been funded by the Deutsche Forschungsgemeinschaft (DFG), under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190.
Appendix A LANCZOS ALGORITHM WITH TEMPORAL DERIVATIVES OF THE AUTOCORRELATION FUNCTION
The construction of the AOTFs in Eq. (1) may be understood as a reformulation of the standard Lanczos scheme Eq. (20) in the basis given by Eq. (21). Rewriting the standard Lanczos step as
$|\tilde{\mathcal{O}}_{n})=b_{n}|\mathcal{O}_{n})=\mathcal{L}|\mathcal{O}_{n-1})-b_{n-1}|\mathcal{O}_{n-2}),$ (40)
we apply $(\mathcal{O}(t)|$ and conjugate to find
$b_{n}(\mathcal{O}_{n}|\mathcal{O}(t))$ $=(\mathcal{O}_{n-1}|\mathcal{L}|\mathcal{O}(t))-b_{n-1}(\mathcal{O}_{n-2}|\mathcal{O}(t)).$ (41)
Using the Heisenberg equation
$\mathcal{L}|\mathcal{O}(t))=|\left[\mathcal{H},\mathcal{O}(t)\right])=-i\partial_{t}|\mathcal{O}(t)),$ (42)
and the gauge (21) we arrive at
$b_{n}i^{n}\varphi_{n}(t)$ $=-i^{n}\partial_{t}\varphi_{n-1}(t)-b_{n-1}i^{n-2}\varphi_{n-2}$ (43)
$\Leftrightarrow b_{n}\varphi_{n}(t)$ $=-\partial_{t}\varphi_{n-1}(t)+b_{n-1}\varphi_{n-2}(t)$ (44)
$=\tilde{\varphi}_{n}(t)$ (45)
which is the first line in the construction scheme given in Eq. (1). Note that the penultimate line is the discrete Schrödinger equation given in Eq. (22). To argue that an appropriate scalar product is given by
$\langle\partial_{t}^{n}C(t),\partial_{t}^{m}C(t)\rangle:=(-1)^{n}\partial_{t}^{n+m}C(t)|_{t=0}$ (46)
several remarks are in order. First note that whenever the integers $n,m$ have different parity the scalar product vanishes because an odd moment is addressed. Moreover as $n,m$ need to have the same parity to yield a nonzero scalar product the above scalar product is symmetric as $(-1)^{n}=(-1)^{m}$. Consider the following:
$C(t)$ $=(\mathcal{O}|e^{i\mathcal{L}t}|\mathcal{O})$ (47)
$\rightarrow\partial_{t}^{n}C(t)$ $=i^{n}(\mathcal{O}|\mathcal{L}^{n}e^{i\mathcal{L}t}|\mathcal{O})$ (48)
$\rightarrow\partial_{t}^{n}C(t)|_{t=0}$ $=i^{n}(\mathcal{O}|\mathcal{L}^{n}|\mathcal{O}).$ (49)
From (49) we may infer that the above scalar product is positive definite as
$\begin{split}\langle\partial_{t}^{n}C(t),\partial_{t}^{n}C(t)\rangle&=(-1)^{n}\partial_{t}^{2n}C(t)|_{t=0}\\ &=(\mathcal{O}|\mathcal{L}^{2n}|\mathcal{O})\\ &=||\mathcal{L}^{n}|\mathcal{O})||\geq 0,\end{split}$ (50)
where we used that $(-1)^{n}=(-1)^{-n}$ for any natural $n$.
We now show that the Lanczos coefficients obtained from $|\tilde{\mathcal{O}}_{n})$ and $\tilde{\varphi}_{n}(t)$ (as appearing in (20) and (1) respectively) coincide, implying the appropriateness of the above scalar product. To this end, consider $|\tilde{\mathcal{O}}_{n})$ expanded in the basis $\{\mathcal{L}^{k}|\mathcal{O})\}_{k=0}^{n}$
$|\tilde{\mathcal{O}}_{n})=\sum_{k=0}^{n}c_{k}\mathcal{L}^{k}|\mathcal{O}),$ (51)
where the $c_{k}$ are real numbers determined by the Lanczos coefficients. From the structure of the Lanczos algorithm all $c_{k}$ where $k$ has a different parity then $n$ are zero. For the respective Lanczos coefficient we may write
$\begin{split}b_{n}^{2}=(\tilde{\mathcal{O}}_{n}|\tilde{\mathcal{O}}_{n})&=\sum_{k,j}c_{j}c_{k}(\mathcal{O}|\mathcal{L}^{j+k}|\mathcal{O})\\ &=\sum_{k,j}c_{j}c_{k}i^{-(k+j)}\partial_{t}^{k+j}C(t)|_{t=0},\end{split}$ (52)
where we made use of Eq. (49). In order to get the respective representation of $\tilde{\varphi}_{n}(t)$ consider
$\begin{split}(\tilde{\mathcal{O}}_{n}|\mathcal{O}(t))&=\sum_{k}c_{k}(\mathcal{O}|\mathcal{L}^{k}|\mathcal{O}(t))=\sum_{k}c_{k}(\mathcal{O}|\mathcal{L}^{k}e^{i\mathcal{L}t}|\mathcal{O})\\ &=\sum_{k}c_{k}i^{-k}\partial_{t}^{k}C(t).\end{split}$ (53)
Via the gauge (21) both pictures are related. Concretely
$\begin{split}\tilde{\varphi}_{n}(t)=b_{n}\varphi_{n}(t)&=i^{-n}(\tilde{\mathcal{O}}_{n}|\mathcal{O}(t))\\ &=i^{-n}\sum_{k}c_{k}i^{-k}\partial_{t}^{k}C(t).\end{split}$ (54)
Note that the $\tilde{\varphi}_{n}(t)$ are real because since $n$ and $k$ need to have the same parity the factors $i^{-(n+k)}$ are never imaginary. For the respective Lanczos coefficient we may write
$\begin{split}\langle\tilde{\varphi}_{n}(t),\tilde{\varphi}_{n}(t)\rangle&=(-1)^{n}\sum_{k,j}c_{k}c_{j}i^{-(k+j)}\langle\partial_{t}^{j}C(t),\partial_{t}^{k}C(t)\rangle\\ &=(-1)^{n}\sum_{k,j}c_{k}c_{j}i^{-(k+j)}(-1)^{k}\partial_{t}^{j+k}C(t)|_{t=0}\\ &=\sum_{k,j}c_{k}c_{j}\underbrace{(-1)^{n+k}}_{+1}i^{-(k+j)}\partial_{t}^{j+k}C(t)|_{t=0}.\end{split}$ (55)
Again noting that $j,k$ need to have the same parity as $n$, eventually we may infer that Eq. (52) and Eq. (55) coincide.
Appendix B NUMERICAL METHOD